OpenStudy (anonymous):

@igreen same concept x-2y=8 2x+y=7

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

Okay, so let's add 2y to both sides in the first equation: x - 2y = 8 x = 2y + 8 Now we can plug in 2y + 8 for 'x' in the 2nd equation: 2x + y = 7 2(2y + 8) + y = 7 Distribute 2 into the parenthesis: 4y + 16 + y = 7 Add 4y + y, what do you get?

OpenStudy (anonymous):

5y?

OpenStudy (anonymous):

Yes, so we have: 5y + 16 = 7 Subtract 16 to both sides.

OpenStudy (anonymous):

-9? that dosen't seem right to me

OpenStudy (anonymous):

Yes, that gives us: 5y = -9

OpenStudy (anonymous):

Divide 5 to both sides.

OpenStudy (anonymous):

What's -9 / 5?

OpenStudy (anonymous):

And yes, this is right.

OpenStudy (anonymous):

1.8

OpenStudy (anonymous):

-1.8

OpenStudy (anonymous):

Yes, that's correct, so that's our y-value. Plug it into any equation: x - 2y = 8 x - 2(-1.8) = 8 Multiply: x + 3.6 = 8 Subtract 3.6 to both sides, what do you get?

OpenStudy (anonymous):

What's 8 - 3.6? @priea

OpenStudy (anonymous):

4.4

OpenStudy (anonymous):

Yes, so our final answer is (4.4, -1.8).

OpenStudy (anonymous):

are u sure? I got that wrong on my end

OpenStudy (anonymous):

Yes, I am 100% sure..last time we added something wrong.

OpenStudy (anonymous):

okay well thanks. ill be sure to ask u for help next time @iGreen.

OpenStudy (anonymous):

No problem, sorry about last time.